OpenStudy (anonymous):

How do you cancel out E or F in this equation using multiplication? 5e+4f=9 4e+5f=9

jimthompson5910 (jim_thompson5910):

notice how adding up the equations gives you 5e+4f=9 4e+5f=9 -------- 9e+9f=18 we want either the e or f to be eliminated, but that doesn't happen here

jimthompson5910 (jim_thompson5910):

which variable do you want to eliminate? e? or f?

jimthompson5910 (jim_thompson5910):

either one works